OpenStudy (anonymous):

CAN SOMEONE PLEASE EXPLAIN HOW TO DO THIS PLEASE?! the equation Y=12841(1.01)^x represents the average cost of an education at a private college after 1990 what is the average cost in 1990 (=0) ? A:12,969 B12,841 C:13,122 D13,099

OpenStudy (phi):

I think they want you to find y when x is 0 we start with the equation y=12841(1.01)^x replace the x with 0 y=12841*(1.01)^0 and simplify remember order of operations: we do exponents first. also remember anything to the 0 power is 1 examples: 3^0= 1 or 10^0 = 1, and so on

OpenStudy (anonymous):

so (1.01)^0 =1?

OpenStudy (anonymous):

@phi

OpenStudy (phi):

yes 1.01 ^ 0 = 1 what is y ?

OpenStudy (anonymous):

it's B right?because 12,841x1 =12,841

OpenStudy (phi):

exactly

OpenStudy (anonymous):

okay thank you!